# Bayesian Structure Learning in Undirected Gaussian Graphical Models: Literature Review with Empirical Comparison

### v2

@@ 5 Case Study: Human Gene Expression
We use the SS algorithm and the BD-A algorithm to perform inference on a gene network between $p=100$ genes using genetic data of $n=60$ unrelated individuals. Several Bayesian structure learning methods are analyzed on this dataset. See, for example, (Mohammadi & Wit, 2015; Li et al., 2019a; Van den Boom et al., 2022a). This dataset can therefore be seen as a benchmark for Bayesian structure learning. We refer to Stranger et al. (2007) and Bhadra & Mallick (2013) for more information on the collection of these data.
@@
Genes are specific sequences of DNA. Gene expression is the process by which a gene produces a protein that influences the functioning of an organism. Some of these proteins directly influence the organism by, for example, initializing a process to break down food. Other proteins, however, only serve to “activate” other genes, which in turn make proteins that activate other genes, and so on. These activation relationships can be shown in a gene network, where every node corresponds to a gene and every edge indicates an activation relationship between two genes. Discovering these gene networks plays a key role in understanding disease susceptibility and therefore, ultimately, in treatment and public health (Stranger et al., 2007). We obtain a sample $\mathbf{X}$ of our variables $X_{i}$, $i=1,\cdots,p$, where $\mathbf{X}$ is an $n\times p$ matrix with elements $x_{li}$ denoting the level of protein corresponding to gene $i=1,\cdots,p$ measured in individual $l=1,\cdots,n$.
Structure learning depends on the assumption that the data comes from a multivariate normal distribution with a mean zero. However, in the gene data set, the univariate histograms of each gene are clearly not normally distributed. Neither are their means zero, see Figure 2.
Figure 2: Univariate histograms of six randomly selected genes in the human gene expression dataset.
@@
In fact, the normality and zero mean assumptions are rarely met in practice. Fortunately, one can transform our non-Gaussian continuous variables $X_{1},\cdots,X_{p}$ to Gaussian variables $Z:=(Z_{1},\cdots,Z_{p})$ with mean zero, i.e. $Z\sim\mathcal{N}(\mathbf{0},\mathbf{\Sigma}_{Z})$, such that the sparsity of $\mathbf{K}_{Z}=\mathbf{\Sigma}_{Z}^{-1}$ still encodes the conditional (in)dependence between the variables $X_{1},\cdots,X_{p}$ as in (2), see for example Liu et al. (2009, 2012). The huge.npn() function in the Huge R package and the bdgraph.npn() function in the BDgraph R package perform this transformation efficiently. One can now apply Bayesian structure learning algorithms on the transformed dataset $\mathbf{Z}$. These transformations are only possible for continuous data. For discrete or binary data, Gaussian copula graphical models can be used to perform structure learning, see Section 6.1 for an overview.
@@
We choose to work with MCMC algorithms to perform inference on the posterior of the gene network. They offer a wide variety of output and are computationally feasible on this 100 node graph. Specifically, we use the SS algorithm and the BD-A algorithm, but the same steps apply to other MCMC algorithms.
@@
First, we select the tuning parameters. For the SS algorithm, we follow Wang (2015) and select $v_{0}=0.02$ and $v_{1}=2$. One could also go for the more time-consuming approach and run the algorithm several times for different parameters, and select the parameters that perform best according to some criterion (Gan et al., 2019). The BD-A algorithm has no tuning parameters.
@@
Second, we need to select a prior for the graph. Generally, the prior is selected using prior knowledge about the sparsity of the true graph. The resulting prior on the graph is then given by (21), where we select the prior sparsity equal to $\delta=0.2$, in line with previous studies on the gene expression dataset; see for example, Van den Boom et al. (2022a) and Li et al. (2019a).
@@
Now, the correct MCMC settings need to be selected. That is, the number of burn-in iterations, the number of MCMC iterations, and the initial graph $G^{(0)}$ of the Markov Chain. To do so, we run the algorithms several times, each time from a different initial graph, and report the number of edges in each Markov chain state. The results are shown in Figure 3.
@@
Figure 3: Amount of edges in the visited graph per MCMC iteration for 10 replications. The density of the initial graph differs per replication, ranging from $0\%$ (empty graph) to $100\%$ (full graph).
The starting position should not be important, since the Markov chain should converge to the posterior distribution regardless of the starting graph. Figure 3 tells us that this indeed happens. Moreover, we can deduce that the first 20 (SS algorithm) and the first 15,000 iterations (BD-A algorithm) can be discarded as burn-in iterations. To be on the safe side, we select burn-in periods of 50 (SS) and 25,000 (BD-A) iterations. Figure 3 also shows that, after the burn-in period, the number of edges of the visited states differs between the SS algorithm (approx. 1,000) and the BD-A algorithm (approx. 1,500). This discrepancy might be caused by the approximations used by both algorithms. Moreover, the SS and BD-A methods use different priors for the precision matrix resulting in different posteriors.
@@
To select the number of MCMC iterations (after burn-in), we plot the edge inclusion probabilities of ten randomly selected links at every state. The results are shown in Figure 4.
@@
We see that convergence happens in less iterations for the SS algorithm (2,000) than for the BD-A algorithm (400,000). This is due to the good mixing of the SS algorithm versus the slow edge-by-edge exploration of the BD-A algorithm. Figure 4 also reveals for the SS algorithm that the final edge inclusion probabilities are all centered around 0.2. It is no coincidence that this is also the value of our prior sparsity. In fact, it turns out that the edge inclusion probabilities of the SS algorithm are very sensitive to the choice of the prior (Figure 5).
Figure 5: Boxplots of the final edge inclusion probabilities of the SS algorithm.
Potential reasons for this prior sensitivity could be the low number of observations ($n=60$) compared to the number of variables ($p=100$), or the fact that the prior sparsity is directly incorporated in the iterative sampling of the graph (Wang, 2015, Equation 12). Due to its prior sensitivity, the SS algorithm fails to distinguish between edges that are likely and unlikely to be in the true graph. It is not suitable for the dataset and therefore left out in the remainder of this section. The BD algorithm is also sensitive to the graph prior but manages in all cases to find both edges with high and low edge inclusion probabilities (Figure 6).
Figure 6: Boxplots of the final edge inclusion probabilities of the BD algorithm.
Figure 7 shows the gene network inferred by the BD-A algorithm. The results show similarities with the existing literature. The five most connected genes of the K-Horseshoe algorithm (Li et al., 2019a) all appear among the most connected genes of our analysis. Moreover, most tree structures of Figure 7 appear exactly in the recovered network given by Bhadra & Mallick (2013, Figure 4).
Figure 7: Gene networks inferred by the BD-A algorithm. Only the 65 edges with the highest edge inclusion probabilities are shown. The width of the edge indicates the edge inclusion probability and varies from narrow (0.58) to thick (1). The color represents the sign of the partial correlation: red (positive) and blue (negative). See supplementary material S4 for the gene names corresponding to the node numbering.
@@ 6 Conclusion and Future Research
In structure learning, Bayesian methods constitute a powerful alternative to frequentist ones by providing inference on the entire posterior. The argument that the speed and simplicity of frequentist methods are superior, is waning as Bayesian methods have improved significantly in both aspects. Bayesian methods that provide accurate solutions to thousand variable problems within mere minutes are now a reality and easily accessible in software packages. This section starts with a short overview of structure learning literature beyond the basic Gaussian case and ends with suggestions for further research.
@@ 6.1 Extension beyond basic GGMs
Beyond Gaussianity assumption: Graphical modeling for non-Gaussian datasets, including types like single nucleotide polymorphisms (SNP), RNA sequencing (RNA-seq), and household income data, is increasingly relevant in data science. Over the past decade, adapting models to handle such non-Gaussian data has been a major research area. For continuous non-Gaussian datasets, like the gene interaction data discussed in Section 5, a two-step method involving a non-parametric data-Gaussianized transformation (Liu et al., 2009) followed by GGM application is effective. This method also extends to multivariate discrete data, as shown in Jia et al. (2017) and (Liang & Jia, 2023, Chapter 5). For multivariate discrete or count data, Gaussian copula graphical models offer an alternative, as detailed in Vinciotti et al. (2022); Dobra & Mohammadi (2017); Dobra & Lenkoski (2011). For mixed data types, Bayesian structure learning methods like those in Dobra & Lenkoski (2011); Mohammadi et al. (2017), that use a latent GGM with a copula approach Hoff (2007) are suitable.
@@
Coloured Gaussian graphical models: these are a class of GGMs with additional symmetry restrictions in the form of equality constraints on the parameters (Højsgaard & Lauritzen, 2008). Most of the methods mentioned in Section 3 can be applied to the coloured GGMs. For example, Li et al. (2020) implemented the RJMCMC method by utilizing the colored G-Wishart prior as the Diaconis-Ylvisaker conjugate prior for GGMs. Roverato & Nguyen (2022) applied a stepwise model search procedure to estimate a brain network from fMRI data.
Time-varying GGMs: While extensive literature revolves around learning static graphical time-invariant models, the change of interdependencies with a covariate (e.g., time or space) is often the rule rather than the exception for real-world data, such as friendships between individuals in a social community, communication between genes in a cell, equity trading between companies, and computer network traffic. Bayesian approaches in time-varying GGMs are computationally burdensome with time complexity. Thus, Yu et al. (2022) developed a low-complexity Bayesian approach by imposing temporally dependent spike and slab priors on the graphs such that they are sparse and vary smoothly across time.
Multivariate regressions: GGMs, and sparse graphical models more broadly, are applicable for statistical inference in high-dimensional regression. This application is straightforward when the response variables and explanatory variables are collectively viewed as a set of p-dimensional variables $X$ within our graphical models, as discussed in Section 6 of Lenkoski & Dobra (2011), for example. In such instances, the inference process can address both the selection of relevant variables and the quantification of uncertainty for the regression coefficients, as outlined in the works of Liang et al. (2022, 2013).
@@ 6.2 Future Research
In the coming decade, Bayesian methods have the potential to become even faster, expand their theoretical guarantees, grow beyond the Gaussian case, and make more impact with their applications. This section conveys how.
First, new MCMC methods can increase the efficiency and feasible dimension of Bayesian structure learning. Mohammadi et al. (2023b) showed the potential of MCMC methods that move only over the graph space, and not over the space of precision matrices. These methods, however, still only allow graphs to change at most one edge per MCMC iteration. It is an open question whether the balance conditions still hold when allowing changes of multiple edges. If true, this could lead to a significant reduction in computation time. Similarly, Van den Boom et al. (2022a) showed the benefit of the informed proposal and delayed acceptance techniques for the reversible jump approach on the joint space of graphs and precision matrices. It remains to be shown whether methods on the graph space can benefit from the same techniques. Wang (2015) introduced the spike and slab prior in Bayesian structure learning. He circumvented the calculation of the normalizing constant by putting its inverse in the prior of the graph. This leads to a computationally efficient method but creates a challenge when incorporating any knowledge of the graph structure in the prior of the graph. This raises two questions. (i) Could the normalizing constant of the spike and slab prior be approximated? (ii) Could G-Wishart methods on the joint space benefit from a similar trick? Furthermore, despite the success of the spike and slab method (Wang, 2015), there are no MCMC methods with different spike and slab priors. Gan et al. (2019) do use a different spike and slab prior, but then design an expectation-maximization method instead of an MCMC method.
Second, moving away from the standard MCMC approach offers a promising and unexplored perspective to Bayesian structure learning. Dai et al. (2022) argue that Sequential Monte Carlo (SMC) methods remain under-used in statistics, despite several advantages. Van den Boom et al. (2022b) and Tan et al. (2017) show that the SMC approach works for Bayesian structure learning. However, their SMC methods cannot yet compete in terms of computational efficiency with the MCMC methods. Likewise, Nemeth & Fearnhead (2021) outlines the benefits of the stochastic gradient MCMC (SGMCMC) method, and Tan & Friel (2020) designs an SGMCMC method for exponential random graph models. As of now, however, no SGMCMC structure learning method exists. Designing one could advance Bayesian structure learning.
Third, Bayesian structure learning needs more theoretical results. Although the majority of MCMC methods meet the balance conditions, not one of them provides rates of convergence as in Tierney (1994, Section 3.2). Moreover, among the methods in Table 1, only three provide results on the concentration of the posterior around the true value as in Sagar et al. (2024), and only one (Van den Boom et al., 2022b) provides an unbiased estimate. In the future, more theoretical results are essential to the credibility of Bayesian structure learning.
@@
Lastly, the increase in the feasible dimension of Bayesian methods will enhance their applications. In Section 5 we reduce the data set containing several thousands of genes to a mere one hundred genes to make the data feasible for computation, potentially losing valuable information. This kind of dimension reduction will be decreasingly necessary as Bayesian methods improve. Likewise, in neuroscience, the models of the brain no longer have to be simplified to just $100$ areas as in Dyrba et al. (2020). This could potentially improve the understanding of cognitive diseases like Alzheimer’s. The enhancements in Bayesian structure learning also open up new applications. Especially exciting but yet unexplored examples are graph neural networks and large language models, which make use of the dependency networks among a large number of variables.

### v3

@@ 5 Case Study: Human Gene Expression
We use the SS-O algorithm and the BD-A algorithm to perform inference on a gene network between $p=100$ genes using genetic data of $n=60$ unrelated individuals. Several Bayesian structure learning methods are analyzed on this dataset; see, for example, (Mohammadi & Wit, 2015; Li et al., 2019a; Van den Boom et al., 2022a; Mohammadi et al., 2023b). This dataset can therefore be seen as a benchmark for Bayesian structure learning. We refer to Stranger et al. (2007) and Bhadra & Mallick (2013) for more information on the collection of the data.
@@
Genes are specific sequences of DNA. Gene expression is the process by which a gene produces a protein that influences the functioning of an organism. Some of these proteins directly influence the organism by, for example, initializing a process to break down food. Other proteins, however, only serve to “activate” other genes, which in turn make proteins that activate other genes, and so on. These activation relationships can be shown in a gene network, where every node corresponds to a gene and every edge indicates an activation relationship between two genes. Discovering these gene networks plays a key role in understanding disease susceptibility and therefore, ultimately, in treatment and public health (Stranger et al., 2007). We obtain a sample $\mathbf{X}$ of our variables $X_{i}$, $i=1,\dots,p$, where $\mathbf{X}$ is an $n\times p$ matrix with elements $x_{li}$ denoting the level of protein corresponding to gene $i=1,\dots,p$ measured in individual $l=1,\dots,n$.
@@
Structure learning depends on the assumption that the data comes from a multivariate normal distribution with a mean zero. However, in the gene data set, the univariate histograms of each gene are clearly not normally distributed. Neither are their means zero, see Figure 2. In fact, the normality and zero mean assumptions are rarely met in practice. Fortunately, one can transform our non-Gaussian continuous variables $X_{1},\dots,X_{p}$ to Gaussian variables $Z:=(Z_{1},\dots,Z_{p})$ with mean zero, i.e., $Z\sim\mathcal{N}(\mathbf{0},\mathbf{\Sigma}_{Z})$, such that the sparsity of $\mathbf{K}_{Z}=\mathbf{\Sigma}_{Z}^{-1}$ still encodes the conditional (in)dependence between the variables $X_{1},\dots,X_{p}$ as in (2); see, for example Liu et al. (2009, 2012). The huge.npn() function in the Huge R package and the bdgraph.npn() function in the BDgraph R package perform this transformation efficiently. One can now apply Bayesian structure learning algorithms on the transformed dataset $\mathbf{Z}$. These transformations are only possible for continuous data. For discrete or binary data, Gaussian copula graphical models can be used to perform structure learning; see Section 6.1 for an overview.
Figure 2: Univariate histograms of three randomly selected genes in the human gene expression dataset.
@@
We choose to work with MCMC algorithms to perform inference on the posterior of the gene network. These algorithms offer a wide variety of output and are computationally feasible on this 100 node graph. Specifically, we use the SS-O algorithm and the BD-A algorithm, but the same steps apply to other MCMC algorithms.
@@
First, we select the tuning parameters. For the SS-O algorithm, we follow Wang (2015) and select $v_{0}=0.02$, $v_{1}=2$ and $\eta=1$. One could also opt for the more time-consuming approach and run the algorithm several times for different parameters, and select the parameters that perform best according to some criterion (Gan et al., 2019). The BD-A algorithm has no tuning parameters.
@@
Second, we need to select a prior for the graph. Generally, the prior is selected using prior knowledge about the sparsity of the true graph. Here, we select the prior given by (20), where we select the prior sparsity equal to $\delta=0.2$, in line with previous studies on the gene expression dataset; see, for example, Van den Boom et al. (2022a) and Li et al. (2019a).
@@
Now, the correct MCMC settings need to be selected. That is, the number of burn-in iterations, the number of MCMC iterations, and the initial graph $G^{(0)}$ of the Markov Chain. The initial graph should not be important, since the Markov chain should converge to the posterior distribution regardless of the starting graph. To confirm that this indeed happens, we run the algorithms several times, each time from a different initial graph, and report the number of edges in each Markov chain state. The resulting plots (Figure 3) indeed show that, regardless of the initial graph, the Markov chain converges to the same neighbourhood. Moreover, we can deduce that approximately the first 10 (SS-O algorithm) and the first 15,000 iterations (BD-A algorithm) can be discarded as burn-in iterations. To be on the safe side, we select burn-in periods of 100 (SS-O) and 25,000 (BD-A) iterations.
@@
Figure 3: Number of edges in the visited graph per MCMC iteration for 10 replications. The density of the initial graph differs per replication, ranging from $0\%$ (empty graph) to $100\%$ (full graph).
@@
To select the number of MCMC iterations (after burn-in), we plot the edge inclusion probabilities of ten randomly selected links at every state. The resulting plots (Figure 4) show that convergence happens in less iterations for the SS-O algorithm (approximately 5,000) than for the BD-A algorithm (approximately 750,000). This is due to the good mixing of the SS-O algorithm versus the slow edge-by-edge exploration of the BD-A algorithm. We set the amount of MCMC iterations accordingly: 5,000 for the SS-O algorithm and 750,000 for the BD-A algorithm.
@@
Figure 5 shows the gene networks inferred by the BD-A algorithm and the SS-O algorithm. The networks show similarities in the both the sign (positive or negative) and the size of the partial correlations. Moreover, the networks show similarities with the existing literature. The five most connected genes of the K-Horseshoe algorithm (Li et al., 2019a) all appear among the most connected genes of our analysis. Moreover, most tree structures of Figure 5 appear exactly in the recovered network given by Bhadra & Mallick (2013, Figure 4).
Figure 5: Gene networks inferred by the SS-O algorithm (left) and the BD-A algorithm (right). Only the 60 edges with the highest edge inclusion probabilities are shown. The width of the edge indicates the strength of the partial correlation. The color represents the sign of the partial correlation: red (negative) and blue (positive). See supplementary material S5 for the gene names corresponding to the node numbering.
The impact of the prior density on the posterior remains relatively unexplored in Bayesian structure learning literature. In order to test the sensitivity of the algorithms to the priors, we ran the algorithms with a sparse ($1\%$ density), average ($20\%$ density) and dense ($50\%$ density) graph prior. We compare the boxplots of the resulting edge inclusion probabilities in Figure S1 (SS-O algorithm) and in Figure S2 (BD-A algorithm) in Section S4 in the supplementary materials. As expected, a higher prior density results in higher edge inclusion probabilities with the SS-O algorithm being slightly more sensitive to a higher prior density. For all prior densities, however, both algorithms are able to distinguish between edges with a low and high edge inclusion probability.
@@ 6 Conclusion and Future Research
In structure learning, Bayesian methods constitute a powerful alternative to frequentist ones by providing inference on the entire posterior. The argument that the speed and simplicity of frequentist methods are superior is waning as Bayesian methods have improved significantly in both aspects. Bayesian methods that provide accurate solutions to thousand variable problems within mere minutes are now a reality and easily accessible in software packages. This section starts with a short overview of structure learning literature beyond the basic Gaussian case and ends with suggestions for further research.
@@ 6.1 Extension Beyond Basic GGMs
Beyond Gaussianity assumption: Graphical modeling for non-Gaussian datasets, including types like single nucleotide polymorphisms, RNA sequencing, and household income data, is increasingly relevant in data science. Over the past decade, adapting models to handle such non-Gaussian data has been a major research area. For continuous non-Gaussian datasets, like the gene interaction data discussed in Section 5, a two-step method involving a non-parametric data-Gaussianized transformation (Liu et al., 2009) followed by GGM application is effective. This method also extends to multivariate discrete data, as shown in Jia et al. (2017) and Liang & Jia (2024, Chapter 5). For multivariate discrete or count data, Gaussian copula graphical models offer an alternative, as detailed in Vinciotti et al. (2022); Dobra & Mohammadi (2017); Dobra & Lenkoski (2011). For mixed data types, Bayesian structure learning methods like those in Dobra & Lenkoski (2011); Mohammadi et al. (2017), that use a latent GGM with a copula approach Hoff (2007), are suitable.
@@
Coloured Gaussian graphical models: These are a class of GGMs with additional symmetry restrictions in the form of equality constraints on the parameters (Højsgaard & Lauritzen, 2008). Most of the methods mentioned in Section 3 can be applied to the coloured GGMs. For example, Li et al. (2020) implemented an RJ method by utilizing the colored $G$-Wishart prior as the Diaconis-Ylvisaker conjugate prior for GGMs. Roverato & Nguyen (2022) applied a stepwise model search procedure to estimate a brain network from fMRI data.
Time-varying GGMs: While extensive literature revolves around learning static graphical time-invariant models, the change of interdependencies with a covariate (e.g., time or space) is often the rule rather than the exception for real-world data, such as friendships between individuals in a social community, communication between genes in a cell, equity trading between companies, and computer network traffic. Bayesian approaches in time-varying GGMs are computationally burdensome with time complexity. Thus, Yu et al. (2022) developed a low-complexity Bayesian approach by imposing temporally dependent spike-and-slab priors on the graphs such that they are sparse and vary smoothly across time.
Multivariate regressions: GGMs, and sparse graphical models more broadly, are applicable for statistical inference in high-dimensional regression. This application is straightforward when the response variables and explanatory variables are collectively viewed as a set of $p$-dimensional variables within our graphical models, as discussed in Section 6 of Lenkoski & Dobra (2011), for example. In such instances, the inference process can address both the selection of relevant variables and the quantification of uncertainty for the regression coefficients, as outlined in the works of Liang et al. (2022, 2013).
@@ 6.2 Future Research
In the coming decade, Bayesian methods have the potential to become even faster, expand their theoretical guarantees, grow beyond the Gaussian case, and make more impact with their applications. This section speculates how.
First, new MCMC methods can increase the efficiency and feasible dimension of Bayesian structure learning. Mohammadi et al. (2023b) show the potential of MCMC methods that move only over the graph space, and not over the space of precision matrices. These methods, however, still only allow graphs to change at most one edge per MCMC iteration. It is an open question whether the balance conditions still hold when allowing changes of multiple edges. If true, this could lead to a significant reduction in computation time. Similarly, Van den Boom et al. (2022a) show the benefit of the informed proposal and delayed acceptance techniques for the reversible jump approach on the joint space of graphs and precision matrices. It remains to be shown whether methods on the graph space can benefit from the same techniques. Wang (2015) introduces the spike-and-slab prior in Bayesian structure learning. He circumvents the calculation of the normalizing constant by putting its inverse in the prior of the graph. This leads to a computationally efficient method but creates a challenge when incorporating any knowledge of the graph structure in the prior of the graph. This raises two questions: (i) Could the normalizing constant of the spike-and-slab prior be approximated? (ii) Could $G$-Wishart methods on the joint space benefit from a similar trick?
Second, moving away from the standard MCMC approach offers a promising and unexplored perspective to Bayesian structure learning. Dai et al. (2022) argue that Sequential Monte Carlo (SMC) methods remain under-used in statistics, despite several advantages. Van den Boom et al. (2022b) and Tan et al. (2017) show that the SMC approach works for Bayesian structure learning. However, their SMC methods cannot yet compete in terms of computational efficiency with the MCMC methods. Likewise, Nemeth & Fearnhead (2021) outline the benefits of the stochastic gradient MCMC (SGMCMC) method, and Tan & Friel (2020) design an SGMCMC method for exponential random graph models. As of now, however, no SGMCMC structure learning method exists.
Third, Bayesian structure learning needs more theoretical results. Although the majority of MCMC methods meet the balance conditions, not one of them provides rates of convergence as in Tierney (1994, Section 3.2). Moreover, among the methods in Table 1, only one (Van den Boom et al., 2022b) provides an unbiased estimate. In the future, more theoretical results are essential to the credibility of Bayesian structure learning.
@@
Lastly, the increase in the feasible dimension of Bayesian methods will enhance their applications. In Section 5, we reduce the data set containing several thousands of genes to a mere one hundred genes to make the data feasible for computation, potentially losing valuable information. This kind of dimension reduction will be decreasingly necessary as Bayesian methods improve. Likewise, in neuroscience, the models of the brain no longer have to be simplified to just $100$ areas (Dyrba et al., 2020). This could potentially improve the understanding of cognitive diseases like Alzheimer’s. The enhancements in Bayesian structure learning also open up new applications. Especially exciting but yet unexplored examples are graph neural networks and large language models, which make use of the dependency networks among a large number of variables.
Acknowledgements
The authors thank the Editor, Prof. Lee, the Associate Editor, and the two referees for their insightful comments, which have significantly improved this article.
Disclosure
The authors report there are no competing interests to declare.